OpenStudy (jchick):

graph that represents the equation y = 4x - 3

OpenStudy (jchick):

Here is the graph I think that this is the one

OpenStudy (jchick):

@amberxoxo could you tell me if I am correct?

OpenStudy (anonymous):

yes i believe u r

OpenStudy (jchick):

Ok thanks!!

OpenStudy (anonymous):

np

OpenStudy (anonymous):

For a second opinion, yes you are correct y = mx + b m = slope = 4 = 4/1 b = y intercept = -3 (the y-intercept on the graph is also -3)

OpenStudy (jchick):

Ok thanks

OpenStudy (anonymous):

No problem =)